OpenStudy (anonymous):

Identify the x and y intercepts for the equation 4x – 3y = 12.

OpenStudy (amistre64):

given ax+by=c xint = c/a yint = c/b

OpenStudy (anonymous):

Sub x = 0 for y intercept Sub y = 0 for x intercept Solve individually.

OpenStudy (anonymous):

wat jethro said, thats wat i always do

OpenStudy (anonymous):

The line crosses the x-axis when the y coordinate is 0 and the line crosses the y-axis when the x coordinate is zero. This leads to equations \[4x – 3\cdot0 = 12\] and \[4\cdot0 – 3y = 12\] Remember that when asked for intersection points you need x and y coordinates, not just the plain answer you got from the equation.

OpenStudy (anonymous):

tru

OpenStudy (anonymous):

can someone give me the full answer

OpenStudy (anonymous):

\[4x–3⋅0=12\]\[4x=12\]\[x=3\]This is the y intercept. \[4⋅0–3y=12\]\[-3y=12\]\[y=-4\]This is the x intercept. The point where the line crosses the x-axis is (3,0). The point where the line crosses the y-axis is (0,-4).

OpenStudy (amistre64):

you typoed it up; swap out "xint" for "yint"

OpenStudy (anonymous):

Ah, so the y intercept didn't mean the x coordinate where the line crosses the x-axis? Sorry, haven't used this concept in English that much before. :)